Solve the system of linear equations 2x-5y=16 and 4x+3y=6 by using elimination.
Answer by mananth(16946)

2 x -5 y = 16 .............1

4 x + 3 y = 6 .............2
Eliminate y
multiply (1)by 3
Multiply (2) by 5
6 x -15 y = 48
20 x + 15 y = 30
Add the two equations
26 x = 78
/ 26
x = 3
plug value of x in (1)
2 x -5 y = 16
6 -5 y = 16
-5 y = 16 -6
-5 y = 10
y = -2
x=3,y=-2